# Call Me When Necessary: LLMs can Efficiently and Faithfully Reason over Structured Environments

### v1

@@ B.2 Reasoning Path Editing Details
Section 4.4 has covered the basic process of reasoning path editing. Here are some implementation details for preparation step, the purpose is to help LLMs identify the error position of previous reasoning path and edit it according to some tips. For the half-way done instances in the Error Messages, we use “compound node” to indicate all cvt nodes in Freebase  Bollacker et al. (2008), and we just sample some instances of each instantiated relation to showcase the path instances in KG of the previous path. The loss of information from sampling is minor because the messages just inform the LLMs that the relations in the reasoning path can be grounded to some certain instances on KG. For candidates, we use a threshold to constrain the size, if there are too many of them, we filter out using similarity search (same as relation-binding module) according to the original question.
@@
Appendix C Datasets
Statistics of datasets are shown in Table C. We experiment Readi on test sets of all datasets. We adopt Hit@1 and denotation accuracy for KGQA and TableQA, respectively. We evaluate Readi on 3 KGQA and 2 TableQA datasets.
WebQuestionsSP (WebQSP) Yih et al. (2016) contains KGQA questions from Google query logs with up to 2-hop reasoning path on Freebase, mostly requiring a list structured reasoning path.
ComplexWebQuestions (CWQ) Talmor and Berant (2018) is a complex KGQA benchmark, challenging for up to 4-hop reasoning path on Freebase, with 55% multi-constrained questions (a tree structured reasoning path).
MetaQA Zhang et al. (2018) includes KGQA questions from movie domain , with 3 sub-sets for 3 levels of difficulty, denoted as MQA-1H, MQA-2H and MQA-3H.
WikiTableQuestions (WTQ) Pasupat and Liang (2015) contains questions over 421 tables challenging for complex aggregation operations, e.g., count, argmax, and sorting.
WikiSQL Zhong et al. (2017) contains complex question based on Wikipedia tables, requiring comparison, aggregation and arithmetic operations.
@@ Appendix C Datasets
Appendix D Baselines
@@
  • AgentBench Liu et al. (2024b) is an agent-based method asking LLMs to call tools based on tool-discription, history and observations.
@@
  GPT3.5 4.3 2.5 0.64 0.46 0.49
      + editing 3.3 2.8 0.86 0.80 0.22
  ChatGPT4 3.4 2.5 0.74 0.60 0.45
      + editing 3.6 3.2 0.89 0.84 0.25
Table 8: Extensive features of Readi-GPT3.5 and Readi-ChatGPT4’s reasoning path, compared with Golden reasoning path. LPP, LGP, AGP, GSR, IER means length of predict path, length of instantiated path, average instantiation progress, grounding success rate and ending with intermediate nodes rate, respectively.
@@
Appendix E Detailed Features of Reasoning Path
Note that quality of reasoning path is multi-dimensional. To further show some insights of reasoning path by LLMs, we design some metrics, including the instantiation progress, error types, etc. We introduce average Length of predict path (LPP), average length of instantiated path (LIP), average instantiation progress (AIP, average of LIP/LPP for each question), instantiation success rate (ISR) and Compound nodes Ending Rate (CNR), where all length mentioned above refers to number of relations on path. Also, we compare with path by SR Zhang et al. (2022) and RoG Luo et al. (2024) (beam = 1). The Golden Path is obtained by extracting relations from golden logical forms in the dataset, just used for analysis.
The absolute performance of different metrics of Readi are shown in Table 8. Also, Figure 6 demonstrates some instantiate progress of Readi-GPT3.5 and Figure 7 shows some relative results of different metrics, compared with fine-tuned methods and the golden path. Additionally, we show the distribution of number of LLM-calls for reasoning path editing in Figure 8.
It is shown that the initial reasoning path by Readi is already comparable with fine-tuned ones. With some necessary editing, Readi significantly gets closer to golden. In addition, not awared of the environments, a large number of Readi’s initial path (49%) get stuck at compound nodes, while fine-tuned methods well memorize the schemas in structured environments. However, with error messages, Readi substantially alleviates this problems and reaches higher QA performance than fine-tuned methods, again demonstrating that unfaithfulness still exits even with large scale fine-tuning.
@@ Appendix E Detailed Features of Reasoning Path
Table 9: Detailed prompts for modules of Readi.
Appendix F Prompt List
Detailed prompts for each module of Readi is in Table 9. For demonstrations, on CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016) we adopt 6 shots for reasoning path generation and 5 shots for other modules. For MQAZhang et al. (2018), the shot number for reasoning path generation follows Li et al. (2023) and we adopt only 3 shots for editing and reasoning (we donnot design a reasoning module for MQA). For WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017), we use 7, 2, 7 shots for generation, editing and reasoning, respectively.

### v2

@@ B.2 Reasoning Path Editing Details
Section 4.4 covers the basic process of reasoning path editing. Here are some implementation details for the preparation step, the purpose is to adopt pertinent structural information on KG to help LLMs identify the error position of previous reasoning path and edit it. For the half-way done instances in the Error Messages, we use “compound node” to indicate all cvt nodes in Freebase  Bollacker et al. (2008), and we just sample several (as a hyper-parameter) instances of each instantiated relation to showcase the path instances in KG of the previous path. The loss of information from sampling is minor because the messages just inform the LLMs that the relations in the reasoning path can be grounded to certain instances on KG. For candidate relations, we also adopt a threshold to constrain the size, if there are too many of them, we filter out using similarity search (same embedding as relation-binding) according to the original question.
@@
Appendix C Detailed Experimental Setups
C.1 Dataset Statistics
We experiment Readi on test sets of all datasets. We adopt Hit@1 and denotation accuracy for KGQA and TableQA, respectively. We evaluate Readi on 3 KGQA and 2 TableQA datasets (Descriptions in Section 5.1). Statistics of datasets are shown in Table 7. Note that we model all datasets as an information retrieval task, instead of a semantic parsing one.
@@ C.1 Dataset Statistics
C.2 LLM API version
For CWQ and WebQSP, we use gpt-3.5-turbo-16k-0613 and gpt-4-32k for GPT3.5 and GPT4, respectively. For others, we adopt gpt-3.5-turbo-0613 and gpt-4 for GPT3.5 and GPT4, respectively.
C.3 Baselines
@@
  • AgentBench Liu et al. (2024) is an agent-based method asking LLMs to call tools based on tool-discription, history and observations.
@@
Appendix D Detailed Analysis
D.1 Elaboration on Reasoning Path Analysis
Note that quality of reasoning path is multi-dimensional. To further show some insights of reasoning path by LLMs, we meticulously design some metrics. We introduce average Length of predicted path (LPP), average length of instantiated path (LIP), average instantiation progress (AIP, average of LIP/LPP for each question), instantiation success rate (ISR) and Compound nodes ending rate (CER), where all length mentioned above refers to number of relations on path. We compare with fine-tuned SR Zhang et al. (2022) and RoG LUO et al. (2024) (beam=1). The Golden Path is obtained by extracting relations from golden logical forms in the dataset, which is only used for analysis.
@@ D.1 Elaboration on Reasoning Path Analysis
Readi-GPT3.5 - init 4.3 2.5 0.64 0.46 0.49
Readi-GPT3.5 - full 3.3 2.8 0.86 0.80 0.22
Readi-GPT4 - init 3.4 2.5 0.74 0.60 0.45
Readi-GPT3.5 - full 3.6 3.2 0.89 0.84 0.25
Table 8: Extensive features of Readi-GPT3.5 and Readi-GPT4’s reasoning path, compared with Golden reasoning path. LPP, LIP, AIP, ISR, IER means length of predict path, length of instantiated path, average instantiation progress, instantiation success rate and ending with intermediate nodes rate, respectively.
The absolute performance of different metrics of Readi are shown in Table 8. Also, Figure 6 demonstrates some instantiate progress of Readi-GPT3.5. Figure 7 shows some relative results of different metrics for Readi-GPT3.5, compared with fine-tuned methods and the Golden. Additionally, we show the distribution of number of LLM-calls for editing for Readi-GPT3.5 in Figure 8.
@@
It is shown that the initial reasoning path by Readi is already comparable with fine-tuned ones. With some necessary editing, Readi significantly gets closer to golden. In addition, without awareness of the environments, a large number of Readi’s initial path (49%) get stuck at compound nodes, while fine-tuned methods well memorize the schemas in structured environments. However, with pertinent feedback upon instantiation errors, Readi substantially alleviates this problems and reaches higher QA performance than fine-tuned methods, again demonstrating that unfaithfulness still exits even with large-scale fine-tuning.
@@
D.2 Performance of varied MAX_EDIT_TIME
We have reported the distribution of Editing times in Figure 5 (GPT4) and Figure 8 (GPT3.5) to show the efficiency of Readi. To some extent, the distribution also showcases the quality of the Reasoning Path and Effectiveness of Editing module. Note that we set the MAX_EDIT_TIME to eight to test the quality, which can be set smaller to save more resources. To further provide some insights, we experiment on different MAX_EDIT_TIME setting on CWQ subset in Table 9. Results show that with more chances to edit, the reasoning path gets better, further showing the effectiveness of Editing. And four times of editing already achieves an comparable outcome.
MAX_EDIT_TIME 2 4 6 8
Hit@1 51.6 59.2 59.7 59.8
Table 9: Readi-GPT35’s performance for different MAX_EDIT_TIME on CWQ.
D.3 Performance of Retriever
The performance of retrieval may affect the instantiation. We calculate the recall of retriever in Table 10, showing that the performance of retrival is not the bottleneck of Readi. Moreover, our implementation is adapted from KB-BINDER Li et al. (2023), which has already shown effectiveness of binding relations to KG.
GPT35 GPT4
CWQ 84.6 86.4
WebQSP 95.5 91.9
Table 10: Recall (%) of retriever for instantiation.
Error Types Instantiation Error Answer not Covered QA Error False Negative
CWQ-GPT3.5 22% 28% 30% 20%
WebQSP-GPT3.5 6% 32% 32% 30%
CWQ-GPT4 18% 36% 20% 26%
WebQSP-GPT4 20% 26% 22% 32%
Table 11: Qualitative error analysis of Readi.
CWQ WebQSP
avg token (k) 52.9 37.6
avg cost (\$) 0.090 0.064
Table 12: Average token cost of Readi (reasoning path generation and editing.
D.4 Token Cost
As indicated in our title, we discusses how LLMs can efficiently and faithfully interact with structured environments. By efficiency, we focus on the way of interaction, i.e. less LLM invocations (Section 1), to obtain pertinent information on large-scale structured environments for multi-hop reasoning. As in some industrial circumstances, the times of LLM-calls are strictly constrained. However, some might concern the token cost of Readi as a measure of efficiency, though it is not our priority. Here, we report the average token cost per question in Table 12. We utilize Python library tiktoken, for GPT3.5 with six-shot for generation, 5-shot for editing and MAX_EDIT_TIME set to 4 for CWQ and 2 for WebQSP.
Note that it is non-trivial to directly compare the exact token cost with other methods, for the sake of number and format of examples, uncontrollable LLM-API traffic, and even the difficulty of the task. For example, StructGPT Jiang et al. (2023a) claims to incorporate 32-shot examples, but the exact few-shot prompts are not provided. Readi’s token cost can also be modulated by adjusting the MAX_EDIT_TIME. Furthermore, excessive calls or overly challenging tasks may cause the LLM-API to fail in providing the appropriate content or format, exceeding the MAX_TRY_TIMES, which also affects the token cost. That’s also why we analyze the times of LLM-call as efficiency of our interaction framework in the main body of paper.
D.5 Qualitative Error Analysis
We analyze 200 randomly-sampled error cases of CWQ and WebQSP in Table 11. We divide the error into four categories (Section 4.4 and 6.2): i. "Instantiation Error": Even after Editing, the path is still not fully instantiated. ii. "Answer not Covered": Answer not in retrieved knowledge. iii. "QA Error": Even the answer is covered, the QA output is still wrong (e.g., hallucination). iv. "False Negative": For example, the ground truth of ‘‘the two continents Turkey is in’’ is ‘‘Eurasia’’, and the model output ‘‘Europe and Asia’’.
Based on Table 11, first, our reasoning path can be well instantiated and utilized to answer the multi-hop reasoning question. Second, there is still room for improvement of the retrieved knowledge, which we would like to focus in our future work. Third, the hallucination in QA reasoning still exits, with GPT4 performing better than GPT3.5.
Appendix E Generalizability
Note that our focus is not on comparative analysis of various LLMs. We propose Readi to enhance LLMs reasoning over structured environments. Therefore, we utilize LLMs with strong understanding and reasoning capabilities (e.g., GPT3.5) to show their performance with Readi. Here we discuss generalizability of Readi framework.
Generalizability to domain-specific KG. The concept of reasoning path is a structured representation of a multi-hop reasoning process (Section 3). By the intrinsic planning ability in Section 1, we adopt LLMs strong question understanding and reasoning ability to directly generate the reasoning path. Here is an intuition of Readi: we humans can navigate multi-hop reasoning challenges across various domains by recognizing named entities and relations of the task. LLMs, trained on large-scale natural language corpus, may develop these abilities. Therefore, Readi is designed to operate without the need of domain-specific knowledge.
We have experimented on both large-scaled KG (CWQ and WebQSP) and domain-specific KG (MetaQA), where Readi shows impressive performance (Section 5). One concern is generalization to specific domain with massive relations. Here, we illustrate how Readi handles such cases with an example.
Consider a question ‘‘What’s the nationality of Lebron James’’. Assume that the LLM has few domain knowledge and initially generates ‘‘[Lebron James] nationality’’ as the reasoning path, whereas the ground truth path in KG is ‘‘[Lebron James] born_in - city_of’’.
When instantiation (Section 4.3), we bind ‘‘nationality’’ to relations in the KG but find none of them connected to [Lebron James], so we invoke editing. In the error message, we include relations around [Lebron James], which is [born_in, sex, father_of, …], as a hint for Editing (Section 4.4 and Appendix B.2).
With the semantic understanding that ‘‘born_in’’ can relate to "nationality of Lebron James", the LLM can correct previous path. The ablation of Editing (Section 6.1) and analysis of Reasoning Path after Editing (Section 6.2) further demonstrate these. Note that relations around [Lebron James] is limited, compared with those in the whole KG.
The whole idea mirrors human cognitive strategies when conducting web searching for multi-hop reasoning tasks, where we flexibly adjust our sub-conscious plan based on information in browsers.
LLM CWQ WebQSP
Llama-2-70b-chat 49.2 76.9
Table 13: Readi’s performance with other LLM-API on CWQ and WebQSP.
Genalizability to other LLMs. Here, we show Readi’s geralizability to other LLMs. Due to API availability, we test Readi with Llama-2-70b-chat on CWQ and WebQSP samples. The MAX_EDIT_TIME is set to 4 for CWQ and 2 for WebQSP. For fair comparison, we adopt our GPT3.5-based QA reasoning module. Table 13 shows that Readi can generalize to other open-sourced LLMs.
We also discuss Readi’s generalizability to fine-tuned models. RoG (Section 6.2) tunes a Llama 2 to generate a reasoning path, showing the feasibility of generating a path by fine-tuning. We’ve shown that Readi performs well with Editing by LLM to improve the path. One concern is about Editing with a fine-tuned model, depending on annotations, which we would go deeper in future works.
Appendix F Comparison with ToG and DATER
We further compare Readi with ToG Sun et al. (2024) and DATER Ye et al. (2023).
F.1 With ToG
Think-on-Graph (ToG) is an LLM-based KGQA method, which also adopts an LLM to reason the answer in an information retrieval manner. Here, we discuss the differences between our Readi and ToG. As discussed in Related Works (Section 2), ToG adopts an LLM to step-by-step filter out some Knowledge Graph (KG) instances, similar to beam search. Note that there is no reasoning path, i.e., a structural representation of the question by LLMs, in ToG’s methodology.
Interaction Paradigm. Our Readi’s novelty lies in the following aspects: 1) Required Capability. In ToG, the LLM traverses on KG, similar to beam search, to filter out entities or relations by scoring all candidates at each step, which mainly requires discrimination ability. In contrast, our Readi requires the strong understanding and reasoning ability to maintain a structural representation of the entire question. 2) Way to Introduce the Environment. To obtain structural information in large-scale KG, ToG introduces all relations around an entity, or all tail entities around a relation, as candidates for LLMs to score the distribution. Such process can be cumbersome because there are sometimes massive candidates on KG, e.g., locations in France. On the other hand, in Readi, we collect pertinent reasoning log only upon the instantiation errors. Readi fully unleash the understanding ability of LLMs and ease the burden of multi-turn scoring for long candidate lists. 3) Grounding. To obtain instances, in ToG, LLMs select items at each step, restricted by the beam size, hindering it from reasoning questions requiring logical operations (based on a set of instances for aggregation, comparison, etc). Conversely, for Readi, we introduce a novel instantiation module (Section 4.3) to obtain all instances based on constraints in the reasoning path.
Generalizability. Readi can generalize to more circumstances (Appendix E): 1) Supported Environments, ToG focuses on KG. However, Readi is a more general interaction framework for reasoning over structured environments. We showcase the concrete implementation of both KGQA (Section 4) and TableQA (Appendix A). 2) Supported Reasoning Tasks. ToG, whose KG instances are restricted by the beam size, falls short for questions requiring logical operations, e.g., aggregation, comparison. However, Readi, adopting LLMs to maintain a structural representation of the question, is not restricted by beam size, which can cover such reasoning cases by LLMs reasoning with retrieved instances.
Experimental Results. We do not cover the results of ToG in main results, for the sake of fair comparison. The evaluation metric from the official published code of ToG differs from the standard Hit@1. Conversely, Readi’s evaluation strictly follows Tan et al. (2023), which is also adopted by all compared LLM-based baselines. While ToG’s reported results are not reproducible, here we report Readi’s results based on ToG’s metrics in Table 14, where our Readi, with less LLM-calls, still outperforms ToG, overall.
Method CWQ WebQSP
ToG-GPT3.5 57.1 76.2
Readi-GPT3.5 57.9 77.5
ToG-GPT4 67.6 82.6
Readi-GPT4 69.2 82.4
Table 14: Readi’s performance compared with ToG over ToG’s metric.
F.2 With DATER
We have considered DATER but did not directly compare with it for the following reasons.
The base LLMs. DATER adopts Codex Chen et al. (2021) as the backbone LLM, which is different from ours. Codex performs significantly better than GPT3.5 in TableQA (9.1% higher acc on WTQ Liu et al. (2023)). However, Codex is close-soured and does not offer API-calls any more [^3] https://platform.openai.com/docs/deprecations. Therefore, it is not fair to directly compare Readi (based on GPT3.5) with DATER.
The way of reasoning over Tables. Our experiment is to show if our interaction framework, Readi, can adopt LLMs to effectively obtain useful information from structured environments and then reason the answer. Therefore, we compare with methods modeling TableQA as an Information Retrieval task (refer to Section 3 and Section 5.2). However, DATER uses a text2SQL model to obtain facts in table. This involves external tools (the SQL Interpreter) to handle the logical and numerical operations (e.g., aggregation, compare and calculation), which is not fair to compare with.
Appendix G Prompt List
Note that we do not modify prompts for baseline methods. Prompts for the vanilla LLMs is in Table 15, following Sun et al. (2024). For TableQA, the prompts for vanilla LLMs is the same in Table 17. Detailed prompts for each module of Readi is in Table 16 (KGQA) and Table 17 (TableQA).
For number of few-shot demonstrations, on CWQ Talmor and Berant (2018) and WebQSP Yih et al. (2016) we adopt 6 shots for reasoning path generation and 5 shots for other modules. For MQAZhang et al. (2018), the shot number for reasoning path generation follows Li et al. (2023) and we adopt only 3 shots for editing and reasoning (we donnot design a reasoning module for MQA). For WTQ Pasupat and Liang (2015) and WikiSQL Zhong et al. (2017), we use 7, 2, 7 shots for generation, editing and reasoning, respectively.
Instruction Please answer the question:
Demonstration Example
Q: What state is home to the university that is represented in sports by George Washington Colonials men’s basketball?
A: Washington, D.C..
Table 15: Prompts for vanilla LLMs for KGQA in our experiments.
@@ Appendix G Prompt List
Table 16: Detailed KGQA prompts for modules of Readi.
Prompts for reasoning path generation
You should predict the needed header and rows in a table for the question.
Demonstration Example
Question: what was the last year where this team was a part of the usl a-league?
| year | division | league | regular season | playoffs | open cup | avg. attendance |
| 2001 | 2 | USL A-League | 4th, Western | Quarterfinals | Did not qualify | 7,169 |
Thought:
First, according to headers and example rows, I need the years the team is in usl a-league league and return the latest year, so I need headers "year" and "league".
Second, I need to constrain "league" = "usl a-league" to know the years of this team as part of the "usl a-league", so I need "league": ["usl a-league"].
Chosen Headers: ["year", "league"]
Constrains: "league": ["usl a-league"]
Prompts for reasoning path editing
There are some mistakes in your previous header or constrains of a question. Follow the given feedback, fix your mistakes and give the correct header and constrains.
Demonstration Example
Question: what was the last year where this team was a part of the usl a-league?
| year | division | league | regular season | playoffs | open cup | avg. attendance |
| 2001 | 2 | USL A-League | 4th, Western | Quarterfinals | Did not qualify | 7,169 |
Wrong Answer:
Chosen Headers: ["year", "team"]
Constrains: "Team": ["usl a-league"]
Feedback:
1. Header [’team’] not in candidate Headers. You can only choose headers from ["year", …"avg. attendance"].
Thought: First, previously I chose headers "year" and "yeam", but "team" is not in Header list. Following the feedback, I need the team in "league"="usl a-league" , so I need headers "year" and "league".
Second, I need to constrain "league" = "usl a-league".
Chosen Headers: ["year", "league"]
Constrains: "league": ["usl a-league"]
Prompts for QA reasoning
You should output the answer of question based on a table.
Output your answer in the last line as "Answer: [’your answer’]"!
Demonstration Example
Question: what was the last year where this team was a part of the usl a-league?
Table:
Headers: league, year
item 1: (league, usl a-league); (year, 2001)
item 2: (league, usl a-league); (year, 2002)
item 3: (league, usl a-league); (year, 2003)
item 4: (league, usl a-league); (year, 2004)
Thought:
First, I know the years the teams is a part of usl a-league are 2001, 2002, 2003 and 2004 from the items in Table.
Second, I calculate the last year is 2004, so the answer is [’2004’].
Answer: [’2004’]
Table 17: Detailed TableQA prompts for modules of Readi.